OpenStudy (anonymous):

\[\sqrt{288}+10\sqrt{8}-3\sqrt{128}\]

OpenStudy (anonymous):

my text gives the answer \[4\sqrt{8}\] but can't \[\sqrt{8}\] simplify to \[2\sqrt{2}\]

OpenStudy (anonymous):

\[4\sqrt{8}\] \[8\sqrt{2}\]

OpenStudy (anonymous):

\[\sqrt{8}=\sqrt{4\times 2}=\sqrt{4}\times\sqrt{2}=2\sqrt{2}\]